OpenStudy (anonymous):

Consider the tip of each hand of clock.find the linear velocity in milimmeter per second for each hand. . Hours hand which 18 milimeter long

OpenStudy (anonymous):

what answer?

OpenStudy (anonymous):

Btww...Lesson for you: A smart person will never show how smart he/she is.....he will simply prove it in kind way :)

OpenStudy (anonymous):

I don't know what are talking about

OpenStudy (anonymous):

An hour hand goes around once in an hour, or 3600 seconds. Distance around one hour = 2*pi*r = 36pi Linear distance = pi/1000 mm/s